(a) Find the​ t-value such that the area in the right tail is 0.20 with 13 degrees of freedom

(b) Find the​ t-value such that the area in the right tail is 0.25 with 27 degrees of freedom

​(c) Find the​ t-value such that the area left of the​ t-value is 0.02 with 24 degrees of freedom.

d) Find the critical​ t-value that corresponds to 80% confidence. Assume 27 degrees of freedom.
Answer #1

a)

This is right tailed test, for α = 0.2 and df = 13
Critical value of t is 0.87.


b)

This is right tailed test, for α = 0.25 and df = 27
Critical value of t is 0.684.


c)

This is left tailed test, for α = 0.02 and df = 24
Critical value of t is -2.172.


d)

sample size, n = 28
degrees of freedom, df = n - 1 = 27

Given CI level is 80%, hence α = 1 - 0.8 = 0.2
α/2 = 0.2/2 = 0.1, tc = t(α/2, df) = 1.314